# TrioXpert: An automated incident management framework for microservice system

### v1

@@ IV-B RQ1: How does TrioXpert perform in AD, FT, and RCL?
IV-C RQ2: Does each component contribute to TrioXpert?
@@
$\mathcal{A}3$ shows a marked drop across all tasks, indicating that the MoE-based collaborative reasoning is more robust than a monolithic LLM under multimodal input load. Similarly, $\mathcal{A}4$ and $\mathcal{A}5$ degrade due to the absence of critical support mechanisms for resolving expert disagreement and mitigating hallucinations.
IV-D RQ3: Does TrioXpert generates interpretable results?
We adopt a qualitative approach to evaluate interpretability, as most baseline methods do not produce explicit reasoning evidence. Applying quantitative metrics in such a setting would be misleading due to the lack of comparable outputs from baselines. Instead, we conduct a case study using real-world scenarios on the AIOpsArena platform [31], a scenario-oriented evaluation framework that supports practical assessments of AIOps algorithms in microservice environments.
Fig. 5 presents an illustrative example generated during the deployment of TrioXpert on AIOpsArena. As shown, TrioXpert explicitly generates reasoning evidence by characterizing system status from both numerical and textual perspectives. The output reveals that checkoutservice2-0 and checkoutservice-1 continuously generate 503 call errors, directly causing order processing failures and triggering cascading effects across downstream services. Furthermore, the failure propagation path is clearly outlined, with each step grounded in observed anomalies. This structured reasoning process significantly enhances the interpretability of the root cause analysis by making its source and logical chain explicit.
Figure 5: This is a case of TrioXpert. In this case, TrioXpert outputs the predictions along with corresponding reasoning evidence wrapped in special tokens. The evidence provides a logically rigorous reasoning process for the predictions, offering strong interpretability.
To further validate its interpretability, we invited two experienced OCEs to review a set of reasoning chains generated by TrioXpert. The experts confirmed that the structured reasoning evidence greatly improved their understanding of the system status and the logic behind each prediction. They particularly valued how the multi-perspective reasoning helped reconstruct the incident timeline and identify key failure points. Moreover, even in cases where TrioXpert’s predictions are incorrect, the evidence it generates still provides OCEs with valuable insights and essential support for understanding system status and facilitating rapid incident diagnosis.
@@ V Related work
Methods based on single-modal data. The metrics, logs, and traces form the core support for the operation and maintenance of modern microservice systems. In previous research, many scholars have attempted to achieve incident management by conducting in-depth analysis of a specific modality of data. For example, regarding metrics data, some studies construct dependency graphs to depict the relationships between system components during incidents, with representative methods including CloudRanger [32], MS-Rank [33], AutoMAP [3], and MicroCause [34]. For log data, certain approaches ([35, 36, 37]) first perform log parsing and then apply hierarchical agglomerative clustering to cluster incidents. As for trace data, many methods ([38, 39, 13]) typically use the 3-sigma approach to determine whether delays are abnormal for anomaly detection, followed by spectrum-based methods for root cause localization. However, these approaches rely exclusively on one modality and fail to capture cross-modal cues. When anomalies manifest differently across modalities—or are absent from the chosen one—such methods often break down.
@@
Methods based on Large Language Models. With the emergence of powerful natural language understanding and reasoning capabilities in LLMs, researchers have explored their application in the field of incident management. However, existing works have yet to adequately address the key limitations in the current incident management domain. For instance, Oasis[20] is primarily used for automatic incident summary generation, Xpert[21] focuses on generating KQL query languages, and Nissist[22] is dedicated to recommending failure remediation measures. None of these methods directly cover multiple stages of the incident management lifecycle. Furthermore, a series of works ([25, 5, 23]) have analyzed the root cause and failure type, but they rely solely on diagnostic information of incidents, failing to fully leverage multimodal data such as metrics, logs, and traces. More critically, these systems do not address core limitations of LLMs—especially hallucination and context window constraints—which hinder their reliability in high-stakes operational environments.
TABLE IV: Comparison of Methods Capabilities in Multi-Task, Multimodal, and Interpretability Aspects.
Methods Multi-Task Multimodal Interpretability
AD FT RCL Metric Log Trace
TrioXpert
CloudRanger [32]
MS-Rank [33]
AutoMAP [3]
MicroCause [34]
LogCluster [36]
SBLD [35]
Log3C [37]
MicroRank [13]
TraceRank [38]
T-Rank [39]
ART [1]
DiagFusion [2]
Eadro [4]
DeepHunt [40]
RCACopilot [5]
ICL_RCA [25]
COMET [23]
As shown in Table IV, although prior work has made notable progress in incident management using single-modal analysis, feature-level multimodal fusion, or emerging LLM capabilities, none have successfully integrated the three key incident management tasks (AD, FT, RCL) with full utilization of rich multimodal observability data in an interpretable and scalable manner. Most multimodal systems rely on statistical abstractions over textual data (e.g., log template frequency or span latency), discarding rich semantic content essential for deep analysis. Conversely, recent LLM-based efforts either focus on narrow subtasks (e.g., summarization or mitigation) or operate on isolated log sequences, making them prone to hallucination and logic inconsistency. TrioXpert addresses these gaps through a collaborative reasoning architecture that combines modality-specific preprocessing with LLMs structured under a Mixture-of-Experts design. This enables reliable multi-task incident analysis with fine-grained interpretability and minimal information loss across modalities.
@@ VI-A Limitations and Possible Solutions
Despite the effectiveness of TrioXpert, several limitations remain. (1) Generalization of numerical modeling. The first pipeline employs self-supervised learning over metric features, which limits generalizability across evolving service topologies or unseen deployment environments. Frequent retraining is required to adapt, impeding deployment scalability. Incorporating training-free anomaly detectors, such as PatternMatcher [41], may reduce this retraining overhead while preserving diagnostic sensitivity. (2) Expert coordination and robustness. TrioXpert adopts a dispatch-and-aggregation MoE architecture, which assumes that each expert (numerical, textual) produces independently correct outputs. In practice, error propagation can occur due to inconsistent reasoning or incomplete evidence. Future work may introduce an adaptive correction mechanism (e.g., a Judge LLM) that verifies cross-expert consistency and triggers targeted re-analysis when semantic contradictions are detected.
@@ VI-B Threats to Validity
VI-C Effiency and Impact on Model Size
TrioXpert achieves low-latency inference despite integrating LLM-based summarization and collaborative reasoning. Across all test cases, the average end-to-end diagnosis is completed within 15 seconds, ensuring no adverse impact on Time-to-Mitigate (TTM) and satisfying industrial requirements for rapid incident response. Regarding model size, TrioXpert currently uses 7B-parameter LLMs for each expert module. This configuration, combined with carefully engineered prompts, achieves strong performance on both $\mathscr{D}1$ and $\mathscr{D}2$, while effectively mitigating hallucinations and preserving interpretability. Although larger models may offer marginal performance gains, the increased computational overhead and latency are not justified under current task complexity. That said, in future deployments targeting more intricate failure scenarios, selectively scaling up the model size may become necessary. At present, the 7B setup offers a favorable trade-off between performance, interpretability, and operational efficiency.
VI-D Why Not Multimodal LLM?
Although multimodal LLMs (MLLMs) represent a natural integration pathway, we chose not to pursue this direction due to two key limitations: (1) Inadequate reasoning capability. Current MLLMs (e.g., Qwen2.5-Max, GPT-4o) excel in cross-modal alignment (image-text, audio-text) [42, 43, 44], but exhibit weak symbolic or causal reasoning on structured enterprise logs and metrics. For incident management, where structured reasoning over logs, traces, and time-series signals is essential, MLLMs underperform compared to modular LLM pipelines. (2) Context limitations. MLLMs face context bottlenecks: even with token limits in the million range, they struggle to ingest full snapshots of high-frequency service telemetry. In our datasets, a 10-minute trace window frequently exceeds this bound. This makes them ill-suited for high-throughput diagnostic pipelines that demand both scale and granularity.
@@ VII Conclusion
Effective incident management is essential to ensure the reliability of microservice systems. This paper presents TrioXpert, an end-to-end framework that combines modality-specific preprocessing with collaborative reasoning with large language models (LLMs). By designing tailored pipelines for metrics, logs, and traces, TrioXpert captures the system status from both numerical and textual perspectives. It then employs a Mixture-of-Experts (MoE) architecture to jointly perform anomaly detection, failure triage, and root cause localization, yielding high accuracy and strong interpretability. The proposed approach demonstrates that the use of modality-aware representations and structured LLM collaboration enables a scalable and explainable incident diagnosis. Although TrioXpert is designed for microservice environments, its principles may be generalized to other complex systems that require multisource analysis and high-stakes decision support.

### v2

@@ IV-B RQ1: Overall Performance
IV-C RQ2: Ablation Study
@@
$\mathcal{A}3$ shows a marked drop across all tasks, indicating that the multi-expert collaborative reasoning is more robust than a monolithic LLM under multimodal input load. Similarly, $\mathcal{A}4$ and $\mathcal{A}5$ degrade due to the absence of support mechanisms for resolving expert disagreement and mitigating hallucinations.
IV-D RQ3: Case Study
We adopt a qualitative approach to evaluate interpretability, as most baseline methods do not produce explicit reasoning evidence. Applying quantitative metrics in such a setting would be misleading due to the lack of comparable outputs from baselines. Importantly, TrioXpert has been deployed in Lenovo’s production-grade microservice platform, where it serves as a real-time diagnostic assistant for incident management. In traditional operations, identifying the root cause of a complex incident typically requires three experienced OCEs working collaboratively for approximately 2.5 hours, often cycling through five or more hypotheses before locating the true root cause. By contrast, TrioXpert significantly improves both efficiency and accuracy: it requires 26 seconds on average to produce a diagnosis, and identifies the correct root cause within two prediction attempts. This improvement in efficiency and precision directly translates to reduced TTM and operational cost savings in production environments. We illustrate TrioXpert ’s interpretability and practical value through three real-world incident cases from Lenovo’s system.
IV-D1 Case 1: Disk Space Exhaustion
When users reported 500 errors in a Java application, TrioXpert’s Numerical Expert detected 100% disk usage across all service instances, while the Textual Expert identified critical logs: “ERROR FileSystemError: Failed to write file… No space left on device.” The Incident Expert synthesized these perspectives, correctly attributing the issue to disk space exhaustion despite the application’s multi-replica nature, and explicitly linked the temporary file generation failure to the observed symptoms.
IV-D2 Case 2: Goroutine Leak
For an HTTP proxy service with severe latency issues, TrioXpert’s Numerical Expert revealed steadily increasing CPU utilization correlating with response time degradation. The Textual Expert identified the synchronous invocation pattern and absence of coroutine termination. The Incident Expert connected these observations to diagnose a goroutine leak, explaining how unbounded coroutine creation led to resource exhaustion.
IV-D3 Case 3: Proxy Misconfiguration
When a data center worker exhibited widespread timeouts, TrioXpert’s Numerical Expert detected anomalous network patterns specific to the affected data center. The Textual Expert identified the HTTPS_PROXY environment variable pointing to a foreign data center. The Incident Expert reconstructed the faulty request flow, identifying how local requests were routed through a foreign proxy, causing packet loss.
Two experienced Lenovo OCEs carefully reviewed the reasoning chains across all three cases. They confirmed that TrioXpert’s structured, evidence-based reasoning made root causes transparent and traceable, enabling them to quickly verify diagnostic conclusions without repeating manual investigations. This level of interpretability not only increased their trust in the system’s outputs, but also reduced the effort required for root cause analysis in terms of both time and expert involvement, thereby improving diagnostic efficiency and reducing operational burden in production environments.
@@ V Related work
Methods based on single-modal data. The metrics, logs, and traces form the core support for the operation and maintenance of modern microservice systems. In previous research, many scholars have attempted to achieve incident management by conducting in-depth analysis of a specific modality of data. For example, regarding metrics data, some studies construct dependency graphs to depict the relationships between system components during incidents, with representative methods including CloudRanger [32], MS-Rank [33], AutoMAP [3], and MicroCause [34]. For log data, certain approaches [35, 36, 37] first perform log parsing and then apply hierarchical agglomerative clustering to cluster incidents. As for trace data, many methods [38, 39, 13] typically use the 3-sigma approach to determine whether delays are abnormal for anomaly detection, followed by spectrum-based methods for root cause localization. However, these approaches rely exclusively on one modality and fail to capture cross-modal cues. When anomalies manifest differently across modalities—or are absent from the chosen one—such methods often break down.
@@
Methods based on Large Language Models. With the emergence of powerful natural language understanding and reasoning capabilities in LLMs, researchers have explored their application in the field of incident management. However, existing works have yet to adequately address the key limitations in the current incident management domain. For instance, Oasis[20] is primarily used for automatic incident summary generation, Xpert[21] focuses on generating KQL query languages, and Nissist[22] is dedicated to recommending failure remediation measures. None of these methods directly cover multiple stages of the incident management lifecycle. Furthermore, a series of works [25, 5, 23] have analyzed the root cause and failure type, but they rely solely on diagnostic information of incidents, failing to fully leverage multimodal data such as metrics, logs, and traces. More critically, these systems do not address core limitations of LLMs—especially hallucination and context window constraints—which hinder their reliability in high-stakes operational environments.
@@ VI-A Limitations and Possible Solutions
Despite the effectiveness of TrioXpert, several limitations remain. (1) Generalization of numerical modeling. The first pipeline employs self-supervised learning over metric features, which limits generalizability across evolving service topologies or unseen deployment environments. Frequent retraining is required to adapt, impeding deployment scalability. Incorporating training-free anomaly detectors, such as PatternMatcher [41], may reduce this retraining overhead while preserving diagnostic sensitivity. (2) Expert coordination and robustness. TrioXpert adopts a dispatch-and-aggregation multi-expert collaboration architecture, which assumes that each expert (numerical, textual) produces independently correct outputs. In practice, error propagation can occur due to inconsistent reasoning or incomplete evidence. Future work may introduce an adaptive correction mechanism (e.g., a Judge LLM) that verifies cross-expert consistency and triggers targeted re-analysis when semantic contradictions are detected.
@@ VI-B Threats to Validity
VI-C Efficiency and Impact on Model Size
TrioXpert achieves low-latency inference despite integrating LLM-based summarization and collaborative reasoning. Across all test cases, the average end-to-end diagnosis is completed within 15 seconds, ensuring no adverse impact on TTM and satisfying industrial requirements for rapid incident response. Regarding model size, TrioXpert currently uses 7B-parameter LLMs for each expert module. This configuration, combined with carefully engineered prompts, achieves strong performance on both $\mathscr{D}1$ and $\mathscr{D}2$, while effectively mitigating hallucinations and preserving interpretability. Although larger models may offer marginal performance gains, the increased computational overhead and latency are not justified under current task complexity. That said, in future deployments targeting more intricate failure scenarios, scaling up the model size may become necessary. At present, the 7B setup offers a favorable trade-off between performance, interpretability, and operational efficiency.
@@ VII Conclusion
This work introduces TrioXpert, an end-to-end framework for unified incident management in microservice systems through modality-specific preprocessing and collaborative LLM reasoning. By preserving semantic richness while addressing LLM limitations via structured prompts, it simultaneously achieves significant accuracy improvements across AD, FT, and RCL. The approach shows that modality-aware representations and structured LLM collaboration enable scalable, interpretable incident management. While designed for microservices, its principles can generalize to other complex, high-stakes systems requiring multisource analysis.
VIII Acknowledgement
This work is supported by the Advanced Research Project of China (31511010501), the National Natural Science Foundation of China (62272249, 62302244), the Fundamental Research Funds for the Central Universities (XXX-63253249), and the CCF-Lenovo Blue Ocean Research Fund.